5x - 3y = 15; (5, 2). Please help
Answer by funmath(2933)

5x - 3y = 15; (5, 2).
I assume you are being told to see if (x,y)=(5,2) is a solution to the equation. To do this, substitute 5 in for x, and 2 in for y. Simplify the equation, if both sides equal it is a solution, if not, it's not.
5(5)-3(2)=15
25-6=15
19=15
(5,2) is not a solution to 5x-2y=15
Happy Calculating!!!